OpenStudy (anonymous):

what is the square root of 1?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1

OpenStudy (akshay_budhkar):

it is just 1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1.000000000000000000000000000000000000

OpenStudy (vishal_kothari):

\[\pm1\]

OpenStudy (akshay_budhkar):

\[\sqrt{x^2}=\left| x \right|\]

OpenStudy (akshay_budhkar):

-1 is not taken as the answer

OpenStudy (anonymous):

just 1

OpenStudy (vishal_kothari):

-1^2=1

OpenStudy (anonymous):

different from asking x^2=1 what's x

OpenStudy (turingtest):

imran and akshay are right

OpenStudy (anonymous):

it can't be both , otherwise square root function won't be a function

myininaya (myininaya):

\[x^2=1 \] has two solutions x=-1 or x=1 \[\sqrt{x}=1 \] has one solution x=1

myininaya (myininaya):

so \[\sqrt{1}=1\]

OpenStudy (anonymous):

tnx...